OpenStudy (anonymous):

r^2-10r+24=0 solve trinoimial

OpenStudy (anonymous):

make factors can be written as r^2-6r-4r+24=0 r(r-6)-4(r-6)=0 (r-4)(r-6)=0 can you solve this now ?

OpenStudy (anonymous):

yes thanks alot :)

OpenStudy (anonymous):

welcome:)